when two bodies move towards each other the distnc btween thm diminishes by 16 m in 10 s. If the bodies move in the same dirctn thd distnc btwn them inareases by 3m for 5 s. Find their velocity.
1
Expert's answer
2013-02-15T05:51:55-0500
when two bodies move towards each other:
V = V1 + V2 (1)

V1, V2 - their velocities
V - relative velocity
If the bodies move in the same direction:
V = V1 -V2 (2)

V = S/t
S - distance , t - time
(1) V1+V2 = 16/10
(2) V1-V2 = 3/5
(1)+(2): 2V1 = 22/10 = 2.2
V1 = 1.1 m/s
(1)-(2): 2V2 = 10/10 = 1
V2 = 0.5 m/s
Answer: V1 = 1.1 m/s, V2 = 0.5 m/s